# Learnability Window in Gated Recurrent Neural Networks

### v3

@@ 4.2 LSTM: one-step Jacobian and effective learning rates
Because $T_{t}$ is block upper-triangular, the top-right block of $T_{t,\ell}$ is
$\sum_{m=\ell+1}^{t}E_{m}\,\Phi_{m-1:\ell},\qquad\Phi_{a:b}:=\prod_{j=b+1}^{a}F_{j},\;\;\Phi_{b:b}=I.$
Extracting its diagonal gives the neuron-wise rates
$\gamma^{(0)}_{t,\ell}\ :=\ \mathrm{diagvec}\!\left(\sum_{m=\ell+1}^{t}E_{m}\,\Phi_{m-1:\ell}\right)\in\mathbb{R}^{H},\quad\gamma^{(0,q)}_{t,\ell}=\sum_{m=\ell+1}^{t}e_{m,q}\!\prod_{j=\ell+1}^{m-1}f_{j,q},\ \ e_{m,q}=o_{m,q}\big(1-\tanh^{2}(c_{m,q})\big).$ (13)
@@
From (12), the first-order contribution to the top-right block yields
@@
where $\cdot\,]_{h,c}$ denotes the $H\times H$ block mapping $c_{\ell}\mapsto h_{t}$. Off-diagonal entries encode cross-neuron mixing and are thus the gate contribution cannot be isolated.
@@
With global learning rate $\mu>0$, define the per-lag, per-neuron effective learning rates
$\mu_{t,\ell}\ :=\ \mu\left(\gamma^{(0)}_{t,\ell}+\gamma^{(1)}_{t,\ell}\right)\in\mathbb{R}^{H},\qquad\mu^{(q)}_{t,\ell}=\mu\left(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\right)$. (14)
These are the scalar multipliers that couple the BPTT gradient at lag $(t{-}\ell)$ with parameter updates on a per-neuron basis.
@@ A GRU updates
Write $J_{t}=T_{t}+\mathcal{R}_{t}$ with the diagonal part $T_{t}:=I-Z_{t}$ (per-neuron leak/retention) and $\mathcal{R}_{t}$ collecting the recurrent corrections. Then, for $\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j}$,
@@
Diagonal rates and correction.
The diagonal product $\Phi_{t:\ell}$ yields the leak/retention rates
$\gamma^{(0)}_{t,\ell}=\mathrm{diagvec}(\Phi_{t:\ell})\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j,q}).$ (24)
Multiplicative filtering by the reset gate in the candidate pathway motivates two additional diagonal rates
@@
[…]
@@
With global learning rate $\mu>0$, the per-lag, per-neuron effective learning rates aggregate the diagonal rates and the first-order diagonal correction:
@@ 4.4 Considerations in terms of optimization dynamics
The effective learning rates $\mu_{t,\ell}=(\mu^{(1)}_{t,\ell},\ldots,\mu^{(H)}_{t,\ell})$ in (14) and (26) quantify how gates re-weight BPTT contributions at lag $(t{-}\ell)$ on a per-neuron basis, even under a fixed global learning rate $\mu$ in SGD (1). Zeroth-order terms produce diagonal contributions—retention/expression in LSTM, leak/reset/mixed in GRU—while first-order terms introduce anisotropy: the recurrent matrices $U_{\bullet}$ modulated by gate slopes mix coordinates and steer updates into privileged subspaces. As a result, gated RNNs act as implicit multi-rate optimizers: they set heterogeneous, lag-dependent learning rates across neurons and selectively bias update directions. This interpretation of gates as multi-rate optimizers was initially discovered in [30] for the scalar-gated RNN models in Appendix B.
@@ 5.1 Setup and link to BPTT
We begin by isolating the portion of the BPTT gradient that reflects the influence of past states. Fix a lag $\ell\!\geq\!1$ and a time index $t$ with $t-\ell\!\geq\!1$. Let $s_{t}$ denote the recurrent state used by BPTT (Sec. 3); for LSTM, $s_{t}=[h_{t};c_{t}]$, whereas for GRU and scalar-gate RNNs in Appendix B, $s_{t}=h_{t}$.
Recalling the one-step decomposition in Eq. (3), we isolate the contribution of a past state $s_{t-\ell}$ to the current gradient. Using the effective learning rates coming from first-order expansion of the $\ell$-step Jacobian product $\mathcal{M}_{t,t-\ell}$ and decomposing $B_{t-\ell}(\theta)=[B^{(1)}_{t-\ell}\ \cdots\ B^{(H)}_{t-\ell}$] into neuronwise blocks, we obtain the following first-order approximation:
$\delta_{t}^{\top}\mathcal{M}_{t,t-\ell}\,B_{t-\ell}(\theta)\;\approx\;\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\underbrace{\big\langle\delta_{t},\;B^{(q)}_{t-\ell}(\theta)\big\rangle}_{\zeta^{(q)}_{t,\ell}}$, (27)
where $H$ denotes the number of neurons. Here, $\zeta^{(q)}_{t,\ell}$ naturally arises from gradient descent and measures the dot product between the instantaneous loss gradient and the parameter sensitivity associated with neuron $q$ at lag $\ell$, while $\mu^{(q)}_{t,\ell}$ accumulates the multiplicative attenuation/amplification induced by gates over the $\ell$-step path.
This formulation makes explicit that BPTT transports gradients along multiplicative chains of effective learning rate values. Consequently, learnability over time is not determined solely by Jacobian numerical stability but by the combined geometry of gating and the statistics of the transported gradients.
@@ 5.2 Binary detection and matched statistic
We formalize finite-horizon learnability by asking when information about a past state $s_{t-\ell}$ remains statistically present in the stochastic gradient at time $t$. For each neuron $q$, define the alignment variable
$\zeta^{(q)}_{t,\ell}:=\langle\delta_{t},\,B^{(q)}_{t-\ell}(\theta)\rangle,$
which measures how the instantaneous loss gradient $\delta_{t}=\partial E_{t}/\partial s_{t}$ aligns with the parameter–sensitivity direction propagated from lag $\ell$.
@@ 5.5 Theoretical implications
Combining the Fano-derived sample-complexity bound in Eq. (34) with the learnability-window analysis of Sec. 5.4 yields several structural insights into how gating structure, heavy-tailed noise, and finite data interact.
@@ 6.2 Empirical learnability windows and time-scale spectra
We now report results for the three scalar-gate RNNs. To keep figures readable, we group plots by quantity rather than by model: envelopes and scaling laws, learnability windows, and time-scale spectra.
@@ Appendix A Matrix product expansion via the Fréchet derivative formulation
Appendix B Scalar-gated RNNs: One-step Jacobians and effective learning rates
We introduce scalar-gated RNN models previously considered in [30] that we use in the paper: a per-neuron (diagonal) gate, a shared (global) gate, and a constant gate. All follow the same update template:
@@ Appendix F Proofs of Lemmas
Recall that for the scalar-gated models (const/shared/diag) we have $\mu^{(q)}_{t,\ell}=\mu\big(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\big)$, while for the GRU the zeroth-order coefficient decomposes as $\gamma^{(0,q)}_{t,\ell}+\rho^{(0,q)}_{t,\ell}+\eta^{(0,q)}_{t,\ell}$ (see Eq. (26)), and the LSTM case in Eq. (14) is analogous. In all architectures, the diagonal coefficients entering these terms are products of time–local factors in $[0,1]$, formed by gates and bounded activation derivatives.
@@
For each neuron $q$ and each architecture, the zeroth-order contribution can be written as a finite sum of terms of the form
$Z^{(q)}_{r,t,\ell}\;=\;c^{(q)}_{r}\prod_{j=t-\ell+1}^{t}a^{(q)}_{r,j},\qquad c^{(q)}_{r}\geq 0,\quad a^{(q)}_{r,j}\in[0,1],$
where $r$ indexes a finite set of contributions. Thus
$\gamma^{(0,q)}_{t,\ell}+\rho^{(0,q)}_{t,\ell}+\eta^{(0,q)}_{t,\ell}\;=\;\sum_{r}Z^{(q)}_{r,t,\ell}$.
When we increase the lag from $\ell$ to $\ell{+}1$, each product gains one extra factor:
$Z^{(q)}_{r,t,\ell+1}=Z^{(q)}_{r,t,\ell}\,a^{(q)}_{r,t-\ell}$.
Since $a^{(q)}_{r,t-\ell}\in[0,1]$, we obtain
$0\leq Z^{(q)}_{r,t,\ell+1}\leq Z^{(q)}_{r,t,\ell}$,
so $\ell\mapsto Z^{(q)}_{r,t,\ell}$ is nonincreasing and nonnegative for every $r$. Summing over $r$ preserves both properties.
@@
The diagonal part of the first-order correction $\gamma^{(1,q)}_{t,\ell}$ can be written as a finite sum
$\gamma^{(1,q)}_{t,\ell}\;=\;\sum_{r=1}^{R_{q}}c_{r}^{(q)}\prod_{j\in\mathcal{I}_{r,\ell}}a^{(q)}_{r,j},\qquad c_{r}^{(q)}\geq 0,\quad a^{(q)}_{r,j}\in[0,1],$
where each index set $\mathcal{I}_{r,\ell}$ satisfies $\mathcal{I}_{r,\ell}\subseteq\mathcal{I}_{r,\ell+1}$. Hence
$0\leq\gamma^{(1,q)}_{t,\ell+1}\leq\gamma^{(1,q)}_{t,\ell}$.
@@ Conclusion.
For each neuron $q$, the effective learning rate coefficient $\mu^{(q)}_{t,\ell}$ is a positive constant times a sum of nonnegative, nonincreasing terms, hence $\ell\mapsto\mu^{(q)}_{t,\ell}$ is nonincreasing. Finally, $f(\ell)=\|\mu_{t,\ell}\|_{1}=\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}$ inherits monotonicity by summation. ∎

### v4

@@ 4.2 LSTM: one-step Jacobian and effective learning rates
Because $T_{t}$ is block upper-triangular with a zero $h\!\to\!h$ block, the $(h,c)$ block of the transport $T_{t,\ell}:=\prod_{j=\ell+1}^{t}T_{j}$ reduces to a single surviving path in which the signal remains in the cell state until time $t$:
$T_{t,\ell}\,]_{h,c}=E_{t}\,\Phi_{t:\ell},\qquad\Phi_{a:b}:=\prod_{j=b+1}^{a}F_{j},\;\;\Phi_{b:b}=I.$
In particular, the $c\!\to\!h$ transition can occur only at the final time $t$. Extracting the diagonal gives the neuron-wise rates
$\gamma^{(0)}_{t,\ell}:=\mathrm{diagvec}\!\left(E_{t}\,\Phi_{t:\ell}\right)\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=e_{t,q}\!\prod_{j=\ell+1}^{t}f_{j,q},\ \ e_{t,q}=o_{t,q}\big(1-\tanh^{2}(c_{t,q})\big).$ (13)
@@
From the first-order product expansion (12), the contribution of a single recurrent mixing insertion to the top-right block of the transport is
@@
where $\cdot\,]_{h,c}$ denotes the $H\times H$ block mapping $c_{\ell}\mapsto h_{t}$. Unlike the zeroth-order case, the presence of the recurrent correction $R_{p}$ reintroduces $h\!\to\!h$ transport, allowing a single mixing event to occur at any intermediate time $p\in\{\ell{+}1,\dots,t\}$. This term therefore collects the diagonal components of the first-order Fréchet expansion in Appendix A, corresponding to neuron-wise self-couplings induced by a single insertion of the recurrent mixing blocks. Off-diagonal entries encode cross-neuron interactions and are intentionally discarded, as they do not admit a neuron-wise time-scale interpretation.
@@
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates
$\mu_{t,\ell}\ :=\ \mu\big(\gamma^{(0)}_{t,\ell}+\gamma^{(1)}_{t,\ell}\big)\in\mathbb{R}^{H},\qquad\mu^{(q)}_{t,\ell}=\mu\big(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\big)$. (14)
These quantities act as scalar multipliers that modulate the contribution of gradients originating at lag $(t{-}\ell)$ to parameter updates, providing a neuron-wise characterization of the effective learning dynamics induced by the LSTM gates.
@@ A GRU updates
Write the Jacobian as $J_{t}=T_{t}+\mathcal{R}_{t}$, where
$T_{t}:=I-Z_{t}$
is the diagonal retention operator induced by the update gate, and $\mathcal{R}_{t}$ collects all remaining recurrent mixing terms. For the transport $\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j}$, a first-order expansion yields
@@
Zeroth-order contributions.
The diagonal product $\Phi_{t:\ell}$ yields the primary leak/retention envelope
$\gamma^{(0)}_{t,\ell}=\mathrm{diagvec}(\Phi_{t:\ell})\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j,q}),$ (24)
which governs the zeroth-order neuron-wise time scales associated with the update gate.
In addition, the multiplicative action of the reset gate along the candidate pathway induces further diagonal attenuations that are useful to track explicitly. We therefore define the diagonal pathwise envelopes
@@
which capture pure reset attenuation and mixed update-reset attenuation, respectively. These quantities isolate the shrinkage associated with reset-controlled chains and are reported as complementary diagonal envelopes, rather than as additional terms in the strict $T_{t}+\mathcal{R}_{t}$ product expansion.
First-order diagonal correction.
The first-order diagonal correction associated with recurrent mixing is obtained by extracting the diagonal of the first–order term in (23):
@@
This term collects the diagonal components arising from a single insertion of the recurrent mixing operator into an otherwise gate-diagonal transport. As in the LSTM case, off-diagonal components encode cross-neuron interactions and are intentionally discarded in order to preserve a neuron-wise time-scale interpretation.
@@
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates as
@@
These quantities provide a diagonal, neuron-wise summary of how update and reset gates jointly shape the effective learning dynamics across temporal lags.
@@ 4.4 Considerations in terms of optimization dynamics
The effective learning rates $\mu_{t,\ell}=(\mu^{(1)}_{t,\ell},\ldots,\mu^{(H)}_{t,\ell})$ in (14) and (26) quantify how gates re-weight BPTT contributions at lag $(t{-}\ell)$ on a per-neuron basis, even under a fixed global learning rate $\mu$ in SGD (1). Zeroth-order terms produce diagonal contributions—retention/expression in LSTM, leak/reset/mixed in GRU—while first-order terms introduce anisotropy: the recurrent matrices $U_{\bullet}$ modulated by gate slopes mix coordinates and steer updates into privileged subspaces. As a result, gated RNNs act as implicit multi-rate optimizers: they set heterogeneous, lag-dependent learning rates across neurons and selectively bias update directions. This interpretation of gates as multi-rate optimizers was initially discovered in [30] for the gated RNN models in Appendix B.
@@ 5.1 Setup and link to BPTT
We begin by isolating the portion of the BPTT gradient that reflects the influence of past states. Fix a lag $\ell\!\geq\!1$ and a time index $t$ with $t-\ell\!\geq\!1$. Let $s_{t}$ denote the recurrent state used by BPTT (Sec. 3); for LSTM, $s_{t}=[h_{t};c_{t}]$, whereas for GRU and gated RNNs in Appendix B, $s_{t}=h_{t}$.
Recalling the one-step decomposition in Eq. (3), we isolate the contribution of a past state $s_{t-\ell}$ to the current gradient. Define
$g_{t,\ell}(\theta)\;:=\;\delta_{t}^{\top}\mathcal{M}_{t,t-\ell}\,B_{t-\ell}(\theta)\;\in\;\mathbb{R}^{1\times P},$
where $\delta_{t}=\partial\mathcal{L}_{t}/\partial s_{t}$ is the local loss gradient, $\mathcal{M}_{t,t-\ell}$ is the $\ell$-step Jacobian product
$\mathcal{M}_{t,t-\ell}:=\prod_{k=t-\ell+1}^{t}\frac{\partial s_{k}}{\partial s_{k-1}},$
and $B_{t-\ell}(\theta)=\partial s_{t-\ell}/\partial\theta$ is the parameter Jacobian. The vector $g_{t,\ell}(\theta)$ represents the lag-$\ell$ contribution of the state $s_{t-\ell}$ to the BPTT gradient in parameter space.
Following the first-order approximation introduced in Appendix A, we approximate the Jacobian product by a diagonal operator collecting neuronwise effective learning rates:
$\mathcal{M}_{t,t-\ell}\;\approx\;D_{t,\ell}\;:=\;\mathrm{diag}\!\big(\mu^{(1)}_{t,\ell},\dots,\mu^{(H)}_{t,\ell}\big)$,
where $\mu^{(q)}_{t,\ell}\geq 0$ quantifies the cumulative modulation induced by gating along the $\ell$-step path associated with neuron $q$.
Decomposing the parameter Jacobian into neuronwise blocks
$B_{t-\ell}(\theta)=[B^{(1)}_{t-\ell}\ \cdots\ B^{(H)}_{t-\ell}],\qquad B^{(q)}_{t-\ell}\in\mathbb{R}^{1\times P},$
the lag-$\ell$ gradient contribution admits the following vector-valued first-order approximation:
$g_{t,\ell}(\theta)\;\approx\;\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\delta_{t}^{(q)}\,B^{(q)}_{t-\ell}(\theta),$ (27)
where $\delta_{t}^{(q)}$ denotes the $q$-th component of the local loss gradient at time $t$.
Equation (27) makes explicit that the lag-$\ell$ contribution to the parameter gradient is a vector obtained by summing neuronwise parameter sensitivities, each modulated by a corresponding effective learning rate. Thus, BPTT transports gradients along multiplicative chains of effective learning rate values, while the geometric structure of the parameter Jacobian determines how this transported signal is distributed across parameters.
@@ 5.2 Binary detection and matched statistic
We formalize finite-horizon learnability by asking when information about a past state $s_{t-\ell}$ remains statistically present in the stochastic gradient at time $t$. As shown in Sec. 5.1, the lag-$\ell$ contribution of $s_{t-\ell}$ to the BPTT gradient is the vector $g_{t,\ell}(\theta)\in\mathbb{R}^{1\times P}$ defined in Eq. (27).
@@
We therefore compress the vector-valued quantity $g_{t,\ell}(\theta)$ into a scalar by applying a fixed linear readout. Let $w\in\mathbb{R}^{P}$ be a random unit-norm vector, drawn independently of the data and fixed throughout training and across all model instances. This projection defines a one-dimensional observable that can be interpreted as a random but unbiased probe of the gradient geometry.
Applying this readout to $g_{t,\ell}(\theta)$ and using the first-order approximation Eq. (27), we obtain the scalar neuronwise alignment variables
$\zeta^{(q)}_{t,\ell}:=\delta_{t}^{(q)}\,\langle B^{(q)}_{t-\ell}(\theta),\,w\rangle,$ (28)
so that the lag-$\ell$ contribution reduces to the weighted sum $\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\zeta^{(q)}_{t,\ell}$.
Because the same realization of $w$ is used for all models and training runs, differences in detectability across architectures reflect differences in temporal credit assignment and gradient transport, rather than artifacts of the projection itself. Moreover, since $w$ is drawn uniformly from the unit sphere, its distribution is rotationally invariant. As a result, the projection $\langle B^{(q)}_{t-\ell}(\theta),\,w\rangle$ provides an unbiased random probe of the typical magnitude and relative alignment of the neuronwise parameter-sensitivity directions, without privileging any fixed parameter direction.
The scalar variables $\zeta^{(q)}_{t,\ell}$ quantify how the transported gradient components associated with neuron $q$ align, through the parameter Jacobian, with the fixed readout direction $w$. The effective learning rates $\mu^{(q)}_{t,\ell}$ weight these alignment terms and thus govern how strongly each lag influences parameter updates.
@@ 5.5 Theoretical implications
Combining the Fano-derived sample-complexity bound in Eq. (35) with the learnability window analysis of Sec. 5.4 yields several structural insights into how gating structure, heavy-tailed noise, and finite data interact.
@@ 6.2 Empirical learnability windows and time-scale spectra
We now report results for the three gated RNNs in Appendix B. To keep figures readable, we group plots by quantity rather than by model: envelopes and scaling laws, learnability windows, and time-scale spectra.
@@ Appendix A Matrix product expansion via the Fréchet derivative formulation
Appendix B Diagonally gated RNNs: One-step Jacobians and effective learning rates
We introduce diagonally gated RNN models previously considered in [30] that we use throughout the paper: a per-neuron (diagonal) gate, a shared (global) gate, and a constant gate. All follow the same update template:
@@ Appendix F Proofs of Lemmas
Recall that for the diagonally gated models in Appendix B we write $\mu^{(q)}_{t,\ell}=\mu\big(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\big)$, while for the GRU we additionally track the diagonal zeroth-order envelopes $\gamma^{(0,q)}_{t,\ell},\rho^{(0,q)}_{t,\ell},\eta^{(0,q)}_{t,\ell}$ (see Eq. (26)); the LSTM zeroth-order coefficient is given by Eq. (13). In all cases, the zeroth-order diagonal envelopes are products of time-local gate factors and bounded activation derivatives, hence lie in $[0,1]$.
Since the first-order diagonal correction may contain signed diagonal contributions (e.g. through diagonal entries of recurrent matrices), we prove monotonicity for the absolute diagonal envelope used throughout the empirical pipeline:
[…]
@@
For the diagonally gated RNNs in Appendix B, for each neuron $q$ the zeroth-order term has the form
$\gamma^{(0,q)}_{t,\ell}\;=\;\prod_{j=t-\ell+1}^{t}a^{(q)}_{j},\qquad a^{(q)}_{j}\in[0,1],$
so that $\gamma^{(0,q)}_{t,\ell+1}=\gamma^{(0,q)}_{t,\ell}\,a^{(q)}_{t-\ell}\leq\gamma^{(0,q)}_{t,\ell}$. For the GRU, the additional envelopes $\rho^{(0,q)}_{t,\ell}=\prod_{j=t-\ell+1}^{t}r_{j,q}$ and $\eta^{(0,q)}_{t,\ell}=\prod_{j=t-\ell+1}^{t}(1-z_{j,q})\,r_{j,q}$ obey the same multiplicative extension property and are therefore nonincreasing in $\ell$. For the LSTM, Eq. (13) gives the single surviving zeroth-order path
$\gamma^{(0,q)}_{t,\ell}=e_{t,q}\prod_{j=\ell+1}^{t}f_{j,q},\qquad e_{t,q}\in[0,1],\;f_{j,q}\in[0,1],$
which is again nonincreasing as $\ell$ increases.
@@
Fix $q$. The diagonal first-order term is a finite sum of contributions indexed by the insertion time $p$ (and, in general, by a finite set of diagonal factors produced by the architecture at time $p$). Each such contribution has the form
$C^{(q)}_{p,t,\ell}\;=\;b^{(q)}_{p}\,\prod_{j\in\mathcal{I}_{p,\ell}}a^{(q)}_{j},\qquad a^{(q)}_{j}\in[0,1],$
where $b^{(q)}_{p}$ is $\ell$-independent (possibly signed), and where the index sets satisfy $\mathcal{I}_{p,\ell}\subseteq\mathcal{I}_{p,\ell+1}$ (i.e. increasing the lag can only add further multiplicative gate factors). Therefore,
[…]
@@ Conclusion.
By the triangle inequality,
[…]